OpenStudy (anonymous):

HELPPPPP PLZZZZ 4sqrt{x-1}-sqrt{15x-1}=0

OpenStudy (anonymous):

4 times sqrt(x-1)?

OpenStudy (anonymous):

\[4\sqrt{x-1}-\sqrt{15x-1}=0\]

OpenStudy (anonymous):